OpenStudy (anonymous):

Will Give Medal What is the slope-intercept form equation of the line that passes through (1, 3) and (3, 7)? y = −2x + 1 y = −2x − 1 y = 2x + 1 y = 2x − 1

Mehek (mehek14):

find the slope \(\dfrac{7-3}{3-1}=\dfrac{4}{2}=2\) so the slope is 2x

Mehek (mehek14):

now we find the y-intercept \(y=2x+b\\3=2*1+b\\2*1=2\\3-2=1\\b=1\\y=2x+1\)

OpenStudy (anonymous):

Thank You

Mehek (mehek14):

anytime ^_^

OpenStudy (anonymous):

Easy way.. will work just for this question.Substitute x coordinate in equations and chk the y that u get..whichever equation gives correct y is the answer.